OpenStudy (anonymous):

Solve for x: −3|2x + 6| = −12 x = 1 and x = 5 x = −1 and x = −5 x = −9 and x = 3 No solution

OpenStudy (anonymous):

I think I could solve it, but I want to make sure that I am correct.

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

its no solution .

OpenStudy (anonymous):

Well, I want to see the solving steps.

OpenStudy (anonymous):

i agree with @makiadanielle

OpenStudy (anonymous):

Can I see how you got that?

OpenStudy (anonymous):

just plug in the solutions that are given to you and see if they work is all i can tell you

OpenStudy (anonymous):

That's not how you solve it. @amistre64 @jayzdd

OpenStudy (anonymous):

that is when it asks you what x eaquals because it gives you ceratin solutions and all you have to do is plug them in

OpenStudy (anonymous):

No, I'm pretty sure that you have to divide both sides.

OpenStudy (anonymous):

@ParthKohli Could you possibly show me how to do this correctly?

OpenStudy (anonymous):

it says that you have to solve for x and there for you have to find what x equals and it gives you a set of choices for you to choose from so there for all you have to do is plug in

OpenStudy (amistre64):

-3 |2x+6| = -12 -3, times what, equals -12? if we cannot determine that it is 4, we can divide off the -3 all the same so |2x+6| = 4 , the absolute value of what numbers gives us 4? |4| = 4 |-4| = 4 so we have 2 cases 2x+6 = 4 or 2x+6 = -4